What is the value of small g in c.g.s system.
 
 
Please answer the above question as soon as possible.

Grade:11

5 Answers

Susmita
5 years ago
g=9.81m/s2
now 1m=102cm.
So g=9.81×102cm/s2=981 cm/s2
 
Technical Nafis
5 years ago
C.G.S system is called centimeter gram and second system.
 
so,
 
g= [980(cm/s2)] 
 
Please approve it you  get helped.
 
With very warms regards 
Technical
 
Thankyou??
Hari
5 years ago
Here in this system of cgs g denotes for the value of grams.but in mks system k denotes for kilo grams.mks system is also known as so system
Ankit Jangra
5 years ago
Cgs unit means centimetre gram and second
g(acceleration due to gravity)=9.8m/s^2
In CGS= 9.8×100/1×1=98cm/s^2
Ankit Jangra
5 years ago
Cgs means centimetre gram and second 
g(acceleration due to gravity)=9.8m/s^2
9m/s^2 =9×100/1×1cm/s^2=980m/s^2